OpenStudy (anonymous):

The point (6,-10) is on a circle with center (2,-6). Write the standard equation of the circle.

OpenStudy (anonymous):

circle equation with center (a, b) and radius r: \[(x - a)^{2} + (y - b)^{2} = r^{2}\] center (2, -6): \[(x - 2)^{2} + (y +6)^{2} = r^{2}\] passing through (6, -10): \[(6 - 2)^{2} + (-10 +6)^{2} = r^{2}\] \[4^{2} + (-4)^{2} = r^{2}\] \[16 + 16 = r^{2}\] \[32 = r^{2}\] so, the equation of the circle is: \[(x - 2)^{2} + (y +6)^{2} = 32\] \[x^{2} - 2x - 2x + 4 + y^{2} + 6y + 6y + 36 = 32\] \[x^{2} + y^{2} - 4x + 12y + 8 = 0\]